# Re(l)ality:

### v2

@@ 1 Introduction
In theories with gauge symmetries, such as General Relativity (GR), the term observables refers to quantities that remain invariant under transformations that leave the physical content of the theory unchanged. These transformations, known as gauge transformations, represent mathematical re-expressions of the theory’s underlying formalism. This paper explores how such observables are defined and interpreted.
For a theory to achieve empirical validity, its gauge-invariant observables must correspond to measurable quantities, facilitating comparisons between theoretical predictions and experimental observations. This correspondence is a critical criterion for the acceptance of any scientific theory. [^1] Although a theory cannot be entirely confirmed or falsified by experimental evidence, its ability to allow meaningful comparisons with reality is essential to avoid being labeled as ad hoc.
@@
As Bamonti and Gomes (2024b) argue, one way to construct local, complete observables in GR is through dynamically coupled reference frames. Specifically, when a pair of partial observables forms a valid solution to the dynamics of the theory, only the diagonal action of a dynamical symmetry (such as a diffeomorphism in GR) preserves this solution.
@@ 2. II)
NB: of course, a frame explicit object will be also spatiotemporally explicit, in the specific sense enclosed by the relational locality typical of the use of reference frames. To avoid confusion, I have decided to distinguish between frame-explicit and spatiotemporally explicit objects, preserving the term ‘spatiotemporal’ to indicate objects written in terms of manifold points (if it helps, the reader may also interpret a spatiotemporally explicit object as ‘manifold explicit’). [^4] The concept of spacetime lacks a universally agreed definition, and its interpretation varies across different frameworks. It may be understood as: (i) the manifold $\mathcal{M}$; (ii) the combination $(\mathcal{M},g_{ab})$, which consists of the manifold equipped with the metric field representing the gravitational field; or (iii) the gravitational field $g_{ab}$ alone. In interpretations (i) and (ii), the distinction lies in whether $\mathcal{M}$ is considered an ontologically independent entity—a stage upon which dynamical variables act—or whether it is inseparable from the fields. Interpretation (iii), however, treats $\mathcal{M}$ as a purely mathematical construct without ontological status (see Rovelli and Gaul (2000); Rovelli (2006); Rovelli and Vidotto (2015); Einstein et al. (2015)). An additional perspective considers spacetime as an emergent structure, defined in terms of observables in the Dirac (gauge-invariant) sense. These observables correspond to the happenings of events, which specify the ”when” and ”where” of physical phenomena. In this paper I choose the more conservative option and identify space-time with option (ii).
@@ 2 The Bundle Formalism: A Gauge Perspective on Reference Frames
In this section, I use the fiber-bundle formalism to describe reference frames and relational observables. This approach, widely used in foundational studies of gauge theories, is primarily attributed to the work of Gomes (see, e.g., Gomes (2023a, b) for a rigorous treatment). Let $M$ represent the space of models $m$ of the theory. The space $M$ can be described as a principal bundle with $S$ as its structure group and $[M]:=\{[m]\mid m\in M\}$ as its base space, where $[m]$ identifies the equivalence class of models under the transformations of $\mathcal{S}$. In this formalism, selecting a reference frame involves defining a unique section map $\sigma:[m]\to\sigma([m])\in M$, which smoothly maps equivalence classes of models to individual models in the space . This corresponds to choosing a submanifold in the fiber bundle that intersects each fiber $\mathbb{F}_{m}:=\mathrm{pr}^{-1}([m])$ exactly once, with $\mathrm{pr}:m\to[m]$ being the projection map.
@@
produces a set of 10 scalar functions indexed by $I$ and $J$, constructed from the metric tensor and its derivatives. [^9] Viable reference frames $\Re_{g}$ must be locally invertible. In spacetimes with continuous symmetries, such as metrics admitting Killing vectors, this condition may fail, making the scalars linearly dependent. Thus, linear independence of $\{\Re_{g}^{(I)}\}$ is necessary but not sufficient for viability. For example, onw could have zero physical degrees of freedom — that is, functional dependence — and still have a viable reference frame. Its gauge-invariance follows from the chain rule for the transformation of $\Re_{g}$:
@@ 3 The View from Nowhere: Frame representations are perspectives on an equivalence class
The relational character of empirical measurements has been widely acknowledged throughout the literature. Observations rely on the relationships established between the system under investigation and the reference frame within which measurements are conducted. For instance, Anderson emphasises that all measurements fundamentally involve comparisons between different physical systems (Anderson, 1967, p.128). Similarly, Rovelli (Rovelli, 1991, p.298) and Landau (Landau and Lifshitz, 1987, p.1) stress the indispensable role of reference frames in any measurement process. This reliance on reference frames underscores the difficulty of reconciling the abstract, frame-free ontology of the View from Nowhere with practical empirical methodologies. [^16] Additional support for the relational nature of empirical data comes from the ”Unobservability Thesis,” which posits that symmetry-related models of a system are empirically indistinguishable (Wallace, 2022). Similarly, discussions on empirical (in)equivalence argue that explicit inclusion of observers leads to an ‘immanent’ conception of empirical distinctions, where models differ only if field configurations exhibit relational differences (Pooley and Read, 2021).
@@
The philosophical foundations of the View from Nowhere: Sophistication.
@@
Moreover, Dewar argues that, Sophistication, unlike Eliminativism, does not require an intrinsic characterisation of $[M]$ for an ontological commitment to equivalence classes alone. However, Martens and Read (2020) have criticised this method of defining invariant structures — defined as the interpretational approach by Møller-Nielsen (2017) — for failing to ensure a perspicuous understanding of the shared ontology underlying various models.
In response to this critique, Gomes (2024a) defends the use of reference frames (which he terms representational conventions), formalised through a projection operator $f_{\sigma}:M\to M$ on the fiber bundle $M$. This procedure provides a perspicuous, yet choice-dependent, characterisation of $[M]$. The adoption of $f_{\sigma}$ instead of the section map $\sigma:[M]\to M$ aligns with the external approach’s claim that intrinsic parameterisation of elements $[m]\in[M]$ is unnecessary. Furthermore, Gomes’ approach integrates the external, symmetry-first perspective with the motivational approach, shielding Sophistication from accusations of being “cheap” (Martens and Read (2020); Belot (2017)). The motivational approach, as defined by Møller-Nielsen (2017), asserts that two symmetry-related models can only be considered physically equivalent if a perspicuous characterisation of their shared ontology is provided. According to Gomes, the projection operator serves precisely this purpose.
I argue that Gomes’ approach, while insightful, is ultimately insufficient. Even if one characterises each equivalence class by selecting a ’representative’ (a relational gauge-invariant observable), this does not resolve the underlying ontological commitment to the equivalence classes themselves. Designating a representative only provides a single perspective on the broader structure, leaving the full ontology of the theory incompletely addressed.
Loss of information.
@@ 4 The View from Everywhere: Frame representations are all that exist
In Bamonti and Gomes (2024b), a map m called external diffeomorphism is be introduced to relate the two frames $\{\phi_{r}^{(I)}\}$ and $\{\phi_{b}^{(I)}\}$, functioning analogously to a coordinate transformation. By redefining $\phi_{r}^{(I)}:=X^{I}_{r}$ and $\phi_{b}^{(I)}:=X^{I^{\prime}}_{b}$, the gauge-invariant observables $g_{IJ}(\phi_{r}):=g_{IJ}(X^{I}_{r})$ and $g_{IJ}(\phi_{b}):=g_{I^{\prime}J^{\prime}}(X^{I^{\prime}}_{b})$ are connected via the map m, which operates as follows:
@@
Clearly, this represents a passive diffeomorphism transformation, demonstrating how changes in reference frames preserve the underlying physical quantities while altering their representation. [^18] Due to the one-to-one correspondence between active and passive diffeomorphisms, from the active perspective, the relationship between $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ can be interpreted as the external diffeomorphism $\textbf{d}:=\phi_{r}^{-1}\circ\textbf{m}\circ\phi_{r}$, or equivalently $\textbf{d}:=\phi_{b}\circ\textbf{m}^{-1}\circ\phi_{b}^{-1}$. In GR, local gauge-invariant observables are covariant under frame transformations (see Pitts (2022)). [^19] Pitts actually refers to covariance relative to a coordinate system. Here, I extend this notion to encompass reference frames. Note that this does not mean that there are no frame-invariant quantities in general, only that these quantities are not local (see Table 1).
While the map m provides a ’shared vocabulary’ for translating between two distinct reference frame representations, it is important to emphasise that these represent separate and fully-fledged physical situations.. [^20] In (Belot, 2017, p.954)’s terminology, m is a “physical symmetry”—an isomorphism linking solutions that represent distinct “possibilia”—as opposed to a “gauge symmetry,” which relates solutions that cannot be taken to represent distinct physical states. Consequently, they are not two perspectives of a shared, total physical state as proposed in the View from Nowhere. Unlike the View from Nowhere, the View from Everywhere does not require the existence of such a shared state. Instead, each gauge-invariant observable represents an independent and self-contained physical reality. Under this framework, we do not require any invariant, frame-free structure in our ontology. Instead, the focus shifts to a theory of frame-dependent yet gauge-invariant objects.
@@
As previously discussed, adopting Sophistication while supporting a motivational approach and offering a perspicuous characterisation of the ontology presents significant challenges. Among the proposals, Gomes’ motivational approach shows promise but ultimately falls short within the Sophistication framework. However, Gomes’ motivational approach aligns naturally with the View from Everywhere. By eliminating the need for a shared ontology based on equivalence classes, the very construction of the gauge-invariant observables constituting the fundamental ontology inherently provides a perspicuous characterisation of the ontology.
@@
First, the breaking of the diffeomorphic freedom of theory is precluded by the existence of the external diffeomorphism m. This map relates different reference frame choices212121m serves as the analogue of the transition map between sections of a principal bundle, given by $\mathrm{\bf{t}}_{\sigma\sigma^{\prime}}(\varphi)=f_{\sigma^{\prime}}(\varphi)^{-1}f_{\sigma}(\varphi)$ (see Gomes (2024a))., ensuring GR to remain m-covariant (see also Bamonti and Gomes (2024a)).
Second, the notion of Relality implies that all the available information resides within a frame. Consequently, there is no ‘total’ information that could be lost when a specific reference frame is selected.
Before concluding, I wish to highlight another advantage of the ”View from Everywhere” concerning the choice of reference frame. The choice is entirely arbitrary, raising the question of what guides such a choice.
@@
In the context of the View from Nowhere, reference frame selection is typically driven by pragmatic considerations, aimed at providing a convenient description of an objective, overarching physical reality. These motivations are largely ’conventional,’ reflecting a preference for ease of description rather than any deeper ontological commitment. In fact, the choice of reference frame, by definition, merely offers a conventional, perspectival description. [^22] A contrasting argument is discussed in (Gomes and Butterfield, 2024, sec.3). In the context of electromagnetism, the authors propose that “a choice of gauge need not be a matter of calculational convenience for some specific problem or class of problems, but can be related to a physically natural and general splitting of the electric field.” However, the authors also note that a physically defined gauge choice (e.g., the Coulomb gauge, which splits the electric field into radiative and Coulombic parts, with the latter determined by the instantaneous charge distribution) is still “non-mandatory.” It corresponds to a particular choice of electric field decomposition. Thus, ultimately, it remains, in my view, a “convention.”
In contrast, in the context of the View from Everywhere, a choice of reference frame constitutes an empirical choice, rather than a conventional one. This is because we cannot conventionally choose another reference frame that represents the ‘same’, ‘true’ physical situation. A different choice is in all respects a different reality.

### v3

@@ 1 Introduction
In theoretical frameworks known as gauge theories, such as General Relativity (GR), the concept of observables refers to quantities that remain unchanged under specific transformations called gauge transformations (Dirac, 1964). These are transformations that do not alter the physical content of the theory, also referred to as the ‘observable content’. Consequently, gauge transformations are often understood as mere mathematical re-expressions of the theory’s formalism and are seen as indicators of a representational redundancy in the theory, where multiple mathematical descriptions correspond to the same physical reality.111‘The ‘gauge argument’ is well known in both physical and philosophical literature (see Berghofer et al. (2023) for a recent overview). I will therefore say no more on the subject, assuming the reader is familiar with the main concepts summarised above. As a final point, it should be pointed out that the presence of gauge symmetries is closely related to the threat of indeterminism. It is precisely to save the theory from this threat that the concept of gauge was introduced in Dirac (1964).
The standard assumption in physics is that a procedure of ‘gauge-fixing’ is necessary to remove the descriptive redundancy, ensuring a unique description of physical configurations through the definition of observable quantities. [^2] It is interesting that the verb ‘fix’ in English can mean two things, both of which are correct in this context: (i) adjusting the formalism to eliminate redundant degrees of freedom; (ii) setting a specific condition (called the gauge condition) from a shortlist of possible conditions, in order to construct a single observable quantity from redundant degrees of freedom. Along these lines, this paper explores how observables are constructed and interpreted in GR.
For a theory to achieve empirical validity, its gauge-invariant observables must correspond to measurable quantities, facilitating comparisons between theoretical predictions and experimental observations. This correspondence is a critical criterion for the acceptance of any scientific theory. [^3] Although a theory cannot be entirely confirmed or falsified by experimental evidence, its ability to allow meaningful comparisons with reality is essential to avoid being labeled as ad hoc (Leplin, 1975).
@@
As Bamonti (2023); Bamonti and Gomes (2024b) argue, one way to construct local, complete observables in GR is through dynamically coupled reference frames. Specifically, when a pair of partial observables forms a valid solution to the dynamics of the theory, only the diagonal action of a dynamical symmetry (such as a diffeomorphism in GR) preserves this solution.
@@ 2. II)
NB: of course, a frame explicit object will be also spatiotemporally explicit, in the specific sense enclosed by the relational locality typical of the use of reference frames. To avoid confusion, I have decided to distinguish between frame-explicit and spatiotemporally explicit objects, preserving the term ‘spatiotemporal’ to indicate objects written in terms of manifold points (if it helps, the reader may also interpret a spatiotemporally explicit object as ‘manifold explicit’). [^6] The concept of spacetime lacks a universally agreed definition, and its interpretation varies across different frameworks. It may be understood as: (i) the manifold $\mathcal{M}$; (ii) the combination $(\mathcal{M},g_{ab})$, which consists of the manifold equipped with the metric field representing the gravitational field; or (iii) the gravitational field $g_{ab}$ alone. In interpretations (i) and (ii), the distinction lies in whether $\mathcal{M}$ is considered an ontologically independent entity—a stage upon which dynamical variables act—or whether it is inseparable from the fields. Interpretation (iii), however, treats $\mathcal{M}$ as a purely mathematical construct without ontological status (see Rovelli and Gaul (2000); Rovelli (2006); Rovelli and Vidotto (2015); Einstein et al. (2015)). An additional perspective considers spacetime as an emergent structure, defined in terms of observables in the Dirac (gauge-invariant) sense. These observables correspond to the happening of events, which specify the ”when” and ”where” of physical phenomena in a relational manner. Accordingly, the ‘where and when’ are consequential to the happening of an event (which is a gauge-invariant observable). So it is the happening of the event that determines where and when it happens and not the event happening at a where and when. In this paper I choose the more conservative option and identify space-time with option (ii).
@@ 2 The Bundle Formalism: A Gauge Perspective on Reference Frames
In this section, I use the fiber-bundle formalism to describe reference frames and relational observables. This approach, widely used in foundational studies of gauge theories (see e.g. Healey (2007); Weatherall (2016)), is primarily attributed to the work of Gomes (see, e.g., Gomes (2023a, b) for a rigorous treatment). Let $M$ represent the space of models $m$ of the theory. The space $M$ can be described as a principal bundle with $S$ as its structure group and $[M]:=\{[m]\mid m\in M\}$ as its base space, where $[m]$ identifies the equivalence class of models under the transformations of $\mathcal{S}$. In this formalism, selecting a reference frame involves defining a unique section map $\sigma:[m]\to\sigma([m])\in M$, which smoothly maps equivalence classes of models to individual models in the space . This corresponds to choosing a submanifold in the fiber bundle that intersects each fiber $\mathbb{F}_{m}:=\mathrm{pr}^{-1}([m])$ exactly once, with $\mathrm{pr}:m\to[m]$ being the projection map.
@@
produces a set of 10 scalar functions indexed by $I$ and $J$, constructed from the metric tensor and its derivatives. [^11] Viable reference frames $\Re_{g}$ must be locally invertible. In spacetimes with continuous symmetries, such as metrics admitting Killing vectors, this condition may fail, making the scalars linearly dependent. Thus, linear independence of $\{\Re_{g}^{(I)}\}$ is necessary for their viability as a reference frame. We should distinguish linear (in)dependence from functional (in)dependence. For example, one could have zero physical degrees of freedom — that is functional dependence — and still have a viable reference frame. Its gauge-invariance follows from the chain rule for the transformation of $\Re_{g}$:
@@ 3 The View from Nowhere: Frame representations are perspectives on an equivalence class
The relational character of empirical measurements has been widely acknowledged throughout the literature. Observations rely on the relationships established between the system under investigation and the reference frame within which measurements are conducted. For instance, Anderson emphasises that all measurements fundamentally involve comparisons between different physical systems (Anderson, 1967, p.128). Similarly, Rovelli (Rovelli, 1991, p.298) and Landau (Landau and Lifshitz, 1987, p.1) stress the indispensable role of reference frames in any measurement process. This reliance on reference frames underscores the difficulty of reconciling the abstract, frame-free ontology of the View from Nowhere with practical empirical methodologies. [^18] Additional support for the relational nature of empirical data comes from the ”Unobservability Thesis,” which posits that symmetry-related models of a system are empirically indistinguishable (Wallace, 2022c). Similarly, discussions on empirical (in)equivalence argue that explicit inclusion of observers leads to an ‘immanent’ conception of empirical distinctions, where models differ only if field configurations exhibit relational differences (Pooley and Read, 2021).
@@
The philosophical foundations of the View from Nowhere: Sophistication
@@
Moreover, Dewar argues that Sophistication, unlike Eliminativism, does not require an intrinsic characterisation of $[M]$ that makes no reference to the elements of $M$ for an ontological commitment to equivalence classes alone. However, Martens and Read (2020) have criticised this method of committing to invariant structures — defined as the interpretational approach by Møller-Nielsen (2017) — for failing to ensure a perspicuous understanding of the shared ontology underlying various models.
The interpretational approach asserts that two symmetry-related models can only be considered physically equivalent if a perspicuous characterisation of their shared ontology is provided.
In contrast, the motivational approach does not require perspicuity as a strict ontological condition but seeks justification for treating symmetry-related models as equivalent.
The ability to provide a perspicuous characterisation of the ontology of the theory aligns more naturally with the interpretational approach, whereas the motivational approach only requires to justify why symmetry-related models can be treated as equivalent.
In response to this critique, Gomes (2024a) defends the use of reference frames (which he terms representational conventions), formalised through a projection operator $f_{\sigma}:M\to M$ on the fiber bundle $M$. This procedure provides a perspicuous, yet choice-dependent, characterisation of $[M]$. The adoption of $f_{\sigma}$ instead of the section map $\sigma:[M]\to M$ aligns with the Sophistication approach’s claim that intrinsic parameterisation of elements $[m]\in[M]$ is unnecessary. Furthermore, Gomes’ approach integrates the interpretational approach with the motivational approach, shielding Sophistication from accusations of being “cheap” (Martens and Read (2020); Belot (2017)) and providing a perspicuous understanding of the ontology. That is, Gomes succeeds in demonstrating that answering why symmetry-related models represent the same physical reality, as required by the motivational approach, also gives us an insightful characterisation of ontology as necessarily required by the interpretational approach.
I argue that Gomes’ approach, while insightful, is ultimately insufficient. Even if one characterises each equivalence class by selecting a ’representative’ (a relational gauge-invariant observable), this does not resolve the underlying ontological commitment to the equivalence classes themselves. Designating a representative only provides a single perspective on the broader structure, leaving the full ontology of the theory incompletely (not perspicuously) addressed.
Loss of information
@@ 4 The View from Everywhere: Frame representations are all that exist
In Bamonti and Gomes (2024b), a map m called external diffeomorphism is introduced to relate the two frames $\{\phi_{r}^{(I)}\}$ and $\{\phi_{b}^{(I)}\}$, functioning analogously to a coordinate transformation. Unlike an ‘ordinary diffeomorphism’, an external diffeomorphism acts directly on the already constructed local, gauge-invariant observables, changing frames and getting us to a different and new observable.
By redefining $\phi_{r}^{(I)}:=X^{I}_{r}$ and $\phi_{b}^{(I)}:=X^{I^{\prime}}_{b}$, the gauge-invariant observables $g_{IJ}(\phi_{r}):=g_{IJ}(X^{I}_{r})$ and $g_{IJ}(\phi_{b}):=g_{I^{\prime}J^{\prime}}(X^{I^{\prime}}_{b})$ are connected via the map m, which operates as follows:
@@
Clearly, this represents a passive diffeomorphism transformation. [^20] Due to the one-to-one correspondence between active and passive diffeomorphisms, from the active perspective, the relationship between $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ can be interpreted as the external diffeomorphism $\textbf{d}:=\phi_{r}^{-1}\circ\textbf{m}\circ\phi_{r}$, or equivalently $\textbf{d}:=\phi_{b}\circ\textbf{m}^{-1}\circ\phi_{b}^{-1}$. This shows that in GR, local gauge-invariant observables are covariant under frame transformations [^21] Pitts (2022) refers to covariance of observables relative to coordinate systems. Here, I propose to extend this notion to encompass reference frames., so the introduction of a gauge choice (in the form of a choice of reference frame) does not spoil either the gauge invariance or the covariance of the theory.
It is important to note that while there can indeed be frame-independent observables, these quantities, however, would inherently be non-local in nature (see Table 1).
While the map m provides a ’shared vocabulary’ for translating between two distinct reference frame representations, it is important to emphasise that these represent separate and fully-fledged physical situations. [^22] In (Belot, 2017, p.954)’s terminology, m is a “physical symmetry”—an isomorphism linking solutions that represent distinct “possibilia”—as opposed to a “gauge symmetry,” which relates solutions that cannot be taken to represent distinct physical states. Consequently, they are not two perspectives of a shared, total physical state as proposed in the View from Nowhere. Unlike the View from Nowhere, the View from Everywhere does not require the existence of such a shared state. Instead, each gauge-invariant observable represents an independent and self-contained physical reality. Under this framework, we do not require any invariant, frame-free structure in our ontology. Instead, the focus shifts to a theory of frame-dependent yet gauge-invariant objects.
@@
As discussed previously, adopting Sophistication while supporting a motivational approach presents significant challenges, particularly when attempting to reconcile it with the demand for a perspicuous characterisation of ontology, which is more closely tied to the interpretational approach. Among the proposals, Gomes’ motivational approach shows promise but ultimately falls short within the Sophistication framework. However, Gomes’ motivational approach aligns naturally with the View from Everywhere. By eliminating the need for a shared ontology based on equivalence classes, the very construction of the gauge-invariant observables constituting the fundamental ontology inherently provides a perspicuous characterisation of the ontology.
@@
First, the breaking of the diffeomorphic freedom of theory is precluded by the existence of the external diffeomorphism m. This map relates different reference frame choices232323m serves as the analogue of the transition map between sections of a principal bundle, given by $\mathrm{\bf{t}}_{\sigma\sigma^{\prime}}(\varphi)=f_{\sigma^{\prime}}(\varphi)^{-1}f_{\sigma}(\varphi)$ (see Gomes (2024a))., ensuring GR to remain m-covariant (see also Bamonti and Gomes (2024a)).
Second, the notion of Relality implies that all the available information resides within each frame. Consequently, there is no ‘total’ information that could be lost when a specific reference frame is selected. This may also alleviate the concerns raised in Wallace (2024) that the choice of certain reference frames (some gauge-fixings) does not preclude the physical features of the system from varying under a gauge transformation, leading to a loss of genuine information and casting doubt on whether they can really be classified as gauge-fixing at all (see ‘The Incompleteness Reason’, ivi, p.13). [^24] To be precise, Wallace studies the so-called ‘unitary gauge’ in electrodynamics. I argue that this choice of gauge corresponds to the use of uncoupled reference frames. In other words, the physical features that change are features of the physical system considered as isolated from the environment (Wallace, 2022a, b). This isolation procedure coincides with the dynamic uncoupling of the reference frame. The “loss of genuine information” referred to by Wallace, therefore, is not really genuine because it is a consequence of this approximate procedure. Within the View from Nowhere, on the other hand, there is the loss of bona-fide genuine information as a result of the use of coupled reference frames, which identify distinct gauge-invariant observables. The View from Nowhere alleviates this, more serious, problem.
@@
Before concluding, I wish to highlight another advantage of the ”View from Everywhere” concerning the choice of reference frame. The choice is entirely arbitrary, raising the question of what guides such a choice. In the context of the View from Nowhere, reference frame selection is typically driven by pragmatic considerations, aimed at providing a convenient description of an objective, overarching physical reality. These motivations are largely ’conventional,’ reflecting a preference for ease of description rather than any deeper ontological commitment. In fact, the choice of reference frame, by definition, merely offers a conventional, perspectival description. [^25] A contrasting argument is discussed in (Gomes and Butterfield, 2024, sec.3). In the context of electromagnetism, the authors propose that “a choice of gauge need not be a matter of calculational convenience for some specific problem or class of problems, but can be related to a physically natural and general splitting of the electric field.” However, the authors also note that a physically defined gauge choice (e.g., the Coulomb gauge, which splits the electric field into radiative and Coulombic parts, with the latter determined by the instantaneous charge distribution) is still “non-mandatory.” It corresponds to a particular choice of electric field decomposition. Thus, ultimately, it remains, in my view, a “convention.” In contrast, in the context of the View from Everywhere, a choice of reference frame constitutes an empirical choice, rather than a conventional one. This is because we cannot conventionally choose one or another reference frame that represents the ‘same’, ‘true’ physical situation in the most helpful way for the purpose at hand. A different choice is in all respects a different reality.
Finally, the View from Everywhere also alleviates the problem whereby the validity of a choice of reference frame cannot be empirically tested. In fact, every measurement is always made within a reference frame, i.e. to make measurements, a reference frame must necessarily be set. Thus, we have no ‘meta-empirical ground’ (that is, a frame-free empirical ground) to compare the various possible reference frames. Within the View from Nowhere, this is a problem, since each choice of reference frame is comparable to a perspective and it makes sense to ask which description is ‘the best’, from a pragmatic or fundamental point of view. The advocate of the View from Everywhere, on the other hand, might claim that there is no need to test the validity of the choice of reference frame, since any choice of reference frame is inherently ’the best’ or ’the most fundamental’, as the only available and existing reality.